A study of identity theft looked at how well consumers protect themselves from this increasingly prevalent crime. The behaviors of 63 college students were compared with the behaviors of 56 nonstudents. One of the questions was "When asked to create a password, I have used either my mother's maiden name, or my pet's name, or my birth date, or the last four digits of my social security number, or a series of consecutive numbers." For the students, 25 agreed with this statement while 28 of the nonstudents agreed.

(a) Display the data in a two-way table.

Students Nonstudents Total
Agreed     1 2 3
Disagreed     4 5 6
Total     7 8 119


Perform the chi-square test. (Round your χ2 to three decimal places and round your P-value to four decimal places.)

χ2 = 9
df = 10
P-value = 11


Summarize the results.

We cannot conclude at the 5% level that students and nonstudents differ in the response to this question. We can conclude at the 5% level that students and nonstudents differ in the response to this question.    


(b) Reanalyze the data using the methods for comparing two proportions that we studied in the previous chapter. Compare the results and verify that the chi-square statistic is the square of the z statistic. (Test students who agreed minus nonstudents who agreed. Round your z to two decimal places and round your P-value to four decimal places.)

z = 13
P-value = 14


(c) The students in this study were junior and senior college students from two sections of a course in Internet marketing at a large northeastern university. The nonstudents were a group of individuals who were recruited to attend commercial focus groups on the West Coast conducted by a lifestyle marketing organization. Discuss how the method of selecting the subjects in this study relates to the conclusions that can be drawn from it.
15
Answer #1

Solution:-

a)

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: The null hypothesis states that the students and non-students does not differ in the response.

Alternate hypothesis: The students and non-students differ in the response.

Formulate an analysis plan. For this analysis, the significance level is 0.05. Using sample data, we will conduct a chi-square test for homogeneity.Alternative hypothesis: At least one of the null hypothesis statements is false.

Analyze sample data. Applying the chi-square test for homogeneity to sample data, we compute the degrees of freedom, the expected frequency counts, and the chi-square test statistic. Based on the chi-square statistic and the degrees of freedom, we determine the P-value.

DF = (r - 1) * (c - 1) = (2 - 1) * (3 - 1)
D.F =
Er,c = (nr * nc) / n

Expected o-EroREdNon-students Expected 25 28.0588210.333456653101 38 34.9411760.267775797187 6363 Students -Erc/ Ec Agreed Disagreed Total 28 24.94117640.3751387347391 28 31.05882350.301247771836 56 53 56 0.6763865065751 119
Er,c
Χ2 = 1.28

where DF is the degrees of freedom, r is the number of populations, c is the number of levels of the categorical variable, nr is the number of observations from population r, nc is the number of observations from level c of the categorical variable, n is the number of observations in the sample, Er,c is the expected frequency count in population r for level c, and Or,c is the observed frequency count in population r for level c.The P-value is the probability that a chi-square statistic having 1 degrees of freedom is more extreme than 1.28.

We use the Chi-Square Distribution Calculator to find P(Χ2 > 1.28) = 0.26.

Interpret results. Since the P-value (0.26) is greater than the significance level (0.05), we have to accept the null hypothesis.

We cannot conclude at the 5% level that students and non-students differ in the response to this question.

b)

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: P1 = P2
Alternative hypothesis: P1eq P2

Note that these hypotheses constitute a two-tailed test. The null hypothesis will be rejected if the proportion from population 1 is too big or if it is too small.

Formulate an analysis plan. For this analysis, the significance level is 0.05. The test method is a two-proportion z-test.

Analyze sample data. Using sample data, we calculate the pooled sample proportion (p) and the standard error (SE). Using those measures, we compute the z-score test statistic (z).

p = (p1 * n1 + p2 * n2) / (n1 + n2)
p = 0.44538
SE = sqrt{ p * ( 1 - p ) * [ (1/n1) + (1/n2) ] }
SE = 0.09128
z = (p1 - p2) / SE

z = - 1.13

where p1 is the sample proportion in sample 1, where p2 is the sample proportion in sample 2, n1 is the size of sample 1, and n2 is the size of sample 2.

Since we have a two-tailed test, the P-value is the probability that the z-score is less than -1.13 or greater than 1.13.

Thus, the P-value = 0.2585

Interpret results. Since the P-value (0.2585) is greater than the significance level (0.05), we have to accept the null hypothesis.

We cannot conclude at the 5% level that students and non-students differ in the response to this question.

Χ2 = 1.28 (From part a)

z = -1.13

z2 = (-1.13)2

z2 = 1.2769

Hence chi ^{2}approx z^{2}